OpenStudy (anonymous):

The first and second year’s sales for a company were $328,000 and $565,000. The expenses for the first year were $117,000. The company had a net profit of $113,000 after two years. What were the company's expenses for the second year? A. $230,000 B. $354,000 C. $663,000 D. $776,000

OpenStudy (pinklion23):

i might be able to help but hold on :)

OpenStudy (anonymous):

@AnswerMyQuestions @KlOwNlOvE @fryfry2003 @Bookworm14 @iSmuggleHuggles @wwhitlock @rainbow_rocks03 @Gemini_Nation_ @wwhitlock

OpenStudy (anonymous):

ok

OpenStudy (pinklion23):

http://openstudy.com/study#/updates/536bb744e4b0e4d5545cf35c this might be right idk

OpenStudy (pinklion23):

oh and if it is pls tell me :)

OpenStudy (rainbow_rocks03):

I think it is C what do you think...IDK though

OpenStudy (anonymous):

can u help with onathor one

OpenStudy (anonymous):

he table shows monthly profits and losses for Drew's fruit stand for a 6-month period. What is the net profit or loss for all six months? Two row table titled Drews Fruit Stand. Data in first row is Jan. Feb. Mar. Apr. May and June. Data in the second row is minus247, 186, minus345, minus127, 322, minus245 A. –$456 B. –$34 C. $456 D. $508

OpenStudy (pinklion23):

hey ways i right for the first question and if i was pls can i have a medal ?! :)

OpenStudy (pinklion23):

i know the answer fir this too LOL

OpenStudy (pinklion23):

i will let them answer the question though so they get a medal too

OpenStudy (pinklion23):

cuz it would be unfair if i get a medal form the whole thing :)

OpenStudy (bookworm14):

It is C add 328,000 + 565,000 =893,000 subtract year 1 expenses : 893,000 - 117,000= 776,000 now subtract your profit: 776,000 - 113,000= 663,000 the total second year expenses= $663,000 Hope This helps! ~Bree

OpenStudy (anonymous):

ok

OpenStudy (pinklion23):

hint: all you need to do is add them all up

OpenStudy (rainbow_rocks03):

it isn't about getting medals @pinklion23 it's about the people you help becoming better in growing knoledge.

OpenStudy (pinklion23):

like so, 1+2+(-4) +7=6

OpenStudy (pinklion23):

well i helped didn't I? @rainbow_rocks03

OpenStudy (pinklion23):

hey@blueribbon8127

OpenStudy (pinklion23):

oops @blueribbon8127 there

OpenStudy (anonymous):

yea

OpenStudy (pinklion23):

you get it?

OpenStudy (anonymous):

im doing the rest of the answers but one is a bit tricky

OpenStudy (anonymous):

for me

OpenStudy (anonymous):

can u help

OpenStudy (pinklion23):

sure just put in a new question

OpenStudy (anonymous):

The table shows Ava's checking account activity for several weeks. Ava's beginning balance was $685. What was her ending balance? Three column table titled Avas Account. Data in the first column is Date, June 12, June 20, June 29, July 2, and July 15. Data in the second column is Deposit, 175, blank, 180, blank, and blank. Data in the third column is Withdrawal, blank, minus240, blank, minus140, and minus350. A. –$45 B. $310 C. $355 D. $995

OpenStudy (pinklion23):

new post sorry

OpenStudy (pinklion23):

oh okay never mind then lol :)

OpenStudy (pinklion23):

what is up with A.

OpenStudy (anonymous):

its -45

OpenStudy (pinklion23):

okay so just add them all up and ta da! you have your answer

OpenStudy (pinklion23):

no 310 is your answer

OpenStudy (anonymous):

oh thanks

OpenStudy (pinklion23):

np

OpenStudy (pinklion23):

hey did some one take my medal? :( well im going to give you a medal cuz you were the one who had to figure it out anyway lol :)

OpenStudy (anonymous):

Use signed numbers to solve. A helicopter descended 1200 feet, rose 800 feet, and then descended 450 feet. What was the net gain or loss in altitude? A. lost 2450 ft B. lost 850 ft C. gained 850 ft D. gained 2450 ft can u check this one i put b

OpenStudy (anonymous):

hehehehhehe

OpenStudy (pinklion23):

okay sure np

OpenStudy (anonymous):

i got a 60 but the ones u did r goood ill just ask my teacher for help

OpenStudy (pinklion23):

okay so -1200+800-450=-850

OpenStudy (anonymous):

but thanks

OpenStudy (pinklion23):

so B. is right good job